On Friday 13/25 of the stocks traded on the NY stock Exchange advanced in price. If 1,495 stocks advanced then how many stocks were traded on that day
Would x bef how many stocks were traded
x=13/25x=1,1495 ????
: On Friday 13/25 of the stocks traded on the NY stock Exchange advanced in price. If 1,495 stocks advanced then how many stocks were traded on that day
Would x bef how many stocks were traded
x=13/25x=1,1495 ????

Answer by solver91311(1877)

13x/25=1495

so:

13x = 1495*25

x=(1495*25)/25

You do the arithmetic.